# The solving degrees for computing Gröbner bases of affine semi-regular polynomial sequences

### v2

@@ 2.3 Extension of the notion of degree of regularity
In this subsection, we shall extend the notion of degree of regularity of a homogeneous ideal $I$ of $R$. Let $\prec$ be a graded order on the monomials of $R$. Recall from Definition 2.1.7 that we set $d_{\rm reg}(I)=\infty$ if $R/I$ is not Artinian, but this is not feasible to analyzing the Gröbner basis computation of $F^{h}$ unless a system defined by $F^{h}$ has no non-trivial root over an algebraic closure. For the feasibility, we here give an alternative definition for degree of regularity, which is viewed as a generalization of the original definition (Definition 2.1.7):
@@
The existence of $d_{0}$ in Definition 2.3.1 is equivalent to that the number of projective zeros of $I$ is finite at most. Note also that, in Definition 2.3.1, we have $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)<\infty$ if $R/I$ is Artinian, and otherwise $\widetilde{d}_{\rm reg}(I)<d_{\rm reg}(I)$ or $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)=\infty$. We also extend the cryptographic semi-regularity (Definition 2.1.9) of a sequence of homogeneous polynomials, as follows:
@@
Here, we relate the solving degree of $F^{h}$ (namely the maximal Gröbner basis degree of $F^{h}$) with our generalized degree of regularity, under some assumptions.
@@
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials. We set $F=\{f_{1},\ldots,f_{m}\}$. Assume that $R/\langle F^{\rm top}\rangle$ is Artinian, and that $\bm{F}^{\rm top}$ is cryptographic semi-regular. Then we have the following:
@@
    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial order $\prec$.
Moreover, the equality holds in the second inequality if $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [36, Section 4] for the original definition).
@@
Put $D:=\widetilde{d}_{\rm reg}(\langle F^{\rm top}\rangle)$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$. By the proof of Theorem 3.1.1 below, we have the following exact sequence:
[…]
@@
for each $d$, where $K_{\bullet}$ denotes the Koszul complex on $\bm{F}^{h}$. Here we have $H_{1}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=0$ and $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}\neq 0$ for $d\leq D-1$, whence $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}<\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}$ for any such $d$. Therefore, the first argument to be proved holds. On the other hand, it follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=0$ for any $d$ with $d\geq D$. Thus, for such any $d$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}$. If $D^{\prime}\geq D$, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1\geq D$. By this together with the surjectivity of $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$, it follows from Lemma 2.2.2 that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ is upper-bounded by $D^{\prime}=\max\{D,D^{\prime}\}$. Also if $D^{\prime}=D-1$, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1=D$. However, $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$ is injective but not surjective, we cannot apply Lemma 2.2.2 for $d_{0}=D^{\prime}$, but apply it for $d_{0}=D^{\prime}+1=D$. Hence, $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ is upper-bounded by $D^{\prime}+1=D=\max\{D,D^{\prime}\}$, as desired.
@@
Assume that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal. First we consider the case where $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, so that $\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$. We put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, and let $\{t_{1},\ldots,t_{r}\}$ be a standard monomial basis of $R^{\prime}_{d}/\langle F^{h}\rangle_{d}$, that is,
$\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle F^{h}\rangle)\rangle}\}.$
@@
For $s\geq 1$, we prove that $\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ is a basis of $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$, from which we have $\dim_{K}R^{\prime}_{d}/\langle F^{h}\rangle_{d}=\dim_{K}R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$ for any positive integer $s$, so that $d\geq\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$ and therefore $d=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$.
@@
By Remark 3.1.2 and Lemma 2.3.5 below, the multiplication-by-$y^{s}$ map from $R_{d}^{\prime}/\langle F^{h}\rangle_{d}$ to $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$ is surjective. Therefore $B_{s}:=\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ generates $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$. Suppose to contrary that $B_{s}$ is not a basis. In this case, there exists an $i$ such that $t_{i}y^{s}$ is divisible by $\mathrm{LM}(g)$ for some $g\in G_{\rm hom}$, but $t_{i}$ is not divisible by $\mathrm{LM}(g)$, where $G_{\rm hom}$ is the reduced Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. Therefore, we can write $t_{i}y^{s}=s_{i}u_{i}y^{s}$ with $s_{i}u_{i}=t_{i}$ and $u_{i}y^{s^{\prime}}=\mathrm{LM}(g)$ for some $s^{\prime}\leq s$. (We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}(g)-\deg u_{i}=s^{\prime}$ and $u_{i}={\rm GCD}(t_{i},{\rm LM}(g))$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}(g)$.
@@
Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$. Then, by the weakly reverse lexicographicness, as ${\rm LM}(g)=u_{i}y^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}(G_{\rm hom})\rangle$. Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}(G_{\rm hom})\rangle$, which is a contradiction.
@@
Finally, we consider the remaining case, namely we have $d_{\rm reg}(\langle F^{\rm top}\rangle)>\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$, so that $\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}=d_{\rm reg}(\langle F^{\rm top}\rangle)$. In this case, it follows from Lemma 2.3.5 below that $d\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, whence $d=d_{\rm reg}(\langle F^{\rm top}\rangle)$. ∎
@@
with $A^{\prime}:=R^{\prime}/\langle F^{h}\rangle$, and the case where $D^{\prime}=D-1$ (i.e., $D^{\prime}+1=D$) means that
@@
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials. We set $F=\{f_{1},\ldots,f_{m}\}$, and assume that $R/\langle F^{\rm top}\rangle$ is Artinian and that $\bm{F}^{\rm top}$ is cryptographic semi-regular. We also suppose that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic. Then we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$.

### v3

@@ 2.3 Extension of the notion of degree of regularity
In this subsection, we shall extend the notion of degree of regularity of a homogeneous ideal $I$ of $R$. Let $\prec$ be a graded ordering on the monomials of $R$. Recall from Definition 2.1.7 that we set $d_{\rm reg}(I)=\infty$ if $R/I$ is not Artinian, but this is not feasible to analyzing the Gröbner basis computation of $F^{h}$ unless a system defined by $F^{h}$ has no non-trivial root over the algebraic closure $\overline{K}$ of $K$. For the feasibility, we here give an alternative definition for degree of regularity, which is viewed as a generalization of the original definition (Definition 2.1.7):
@@
Remark 2.3.2.
For a homogeneous ideal $I$ in $R$, its generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$ is nothing but the index of regularity $i_{\rm reg}(I)$, if $I$ has finitely many projective zeros over the algebraic closure $\overline{K}$. Here, the index of regularity of $I$ is defined as follows: Denoting by ${\rm HP}_{R/I}$ the Hilbert polynomial of $R/I$, we define the index of regularity of $I$ as the smallest non-negative integer $i_{\rm reg}(I)$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for all $d$ with $d\geq i_{\rm reg}(I)$. Note that the index of regularity $i_{\rm reg}(I)$ is also called the Hilbert regularity of $I$, which is often denoted by ${\rm hilb}(I)$. The following three conditions are all equivalent: (1) $I$ has finite number of projective zeros over the algebraic closure $\overline{K}$. (2) ${\rm HP}_{R/I}$ is a constant. (3) $\widetilde{d}_{\rm reg}(I)<\infty$. Therefore, we have $\tilde{d}_{\rm reg}(I)=i_{\rm reg}(I)$. We also remark that ${\rm HS}_{R/I}$ is written as ${\rm HS}_{R/I}(z)=h(z)/(1-z)$ for a unique polynomial $h(z)$, for which $\widetilde{d}_{\rm reg}(I)=\deg(h)$ and $h(1)={\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for all $d$ with $d\geq\widetilde{d}_{\rm reg}(I)=i_{\rm reg}(I)$.
Note that the degree of regularity $d_{\rm reg}(I)$ is also equal to $i_{\rm reg}(I)$ if $R/I$ is Artinian, but these are distinguished in the literature. Following this, we distinguish $\tilde{d}_{\rm reg}(I)$ and $i_{\rm reg}(I)$.
Note also that, in Definition 2.3.1, we have $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)<\infty$ if $R/I$ is Artinian, and otherwise $\widetilde{d}_{\rm reg}(I)<d_{\rm reg}(I)=\infty$ or $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)=\infty$. We also extend the cryptographic semi-regularity (Definition 2.1.10) of a sequence of homogeneous polynomials, as follows:
@@
Here, we relate the solving degree of $F^{h}$ (namely the maximal Gröbner basis degree of $F^{h}$) with our generalized degree of regularity, under some assumptions. For this, we extend the notion of top part to a homogeneous polynomial in $R^{\prime}=R[y]$ as follows:
Definition 2.3.4.
For a homogeneous polynomial $h$ in $R^{\prime}=R[y]$, we call $h|_{y=0}$ the top part of $h$, and denote it by $h^{\rm top}$. For a set $H$ of homogeneous polynomials in $R^{\prime}$, its top part is defined by $H^{\rm top}:=\{h^{\rm top}:h\in F\}\subset R$, and similarly $\bm{H}^{\rm top}$ is defined for a sequence $\bm{H}$ of homogeneous polynomials in $R^{\prime}$.
In Definition 2.3.4, if $h^{\rm top}$ is not zero, it coincides with the top part $(h|_{y=1})^{\rm top}$ of the dehomogenization $h|_{y=1}$ of $h$. We remark that $g^{\rm top}=(g^{h})^{\rm top}$ for a polynomial $g$ in $R$.
First, we prove the following lemma (a generalization of Lemma 2.3.5, see Appendix A.3 below):
Lemma 2.3.5.
Let $H=\{h_{1},\ldots,h_{m}\}$ be a set of homogeneous polynomials in $R^{\prime}=R[y]$. Assume that $R^{\prime}/\langle H,y\rangle$ (which is isomorphic to $R/\langle H^{\rm top}\rangle$ with $H^{\rm top}=H|_{y=0}$) is Artinian, namely $D:=d_{\rm reg}(\langle H^{\rm top}\rangle_{R})<\infty$. We also assume that $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})<\infty$. Then, if $D^{\prime}\geq D$, we have $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\leq D^{\prime}$ for any graded monomial ordering $\prec^{\prime}$ on $R^{\prime}$ given in Lemma 2.2.2.
Proof.
As in the proof of Theorem 3.1.1 below (or considering a mapping cone of Koszul complexes), we have the following exact sequence:
[…]
for each $d$, where $K_{\bullet}$ (resp. $K_{\bullet}^{\prime}$) denotes the Koszul complex on the sequence $(h_{1},\ldots,h_{m})$ (resp. the sequence $(h_{1},\ldots,h_{m},y)$). It follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}^{\prime})_{d}=0$ for any $d$ with $d\geq D$. Thus, for any $d$ with $d\geq D$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle H\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle H\rangle)_{d}$. Here, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1\geq D$. By this together with the surjectivity of $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$ (this surjectivity comes from $D^{\prime}\geq D$), it follows from Lemma 2.2.2 that $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)$ is upper-bounded by $D^{\prime}$, as desired. ∎
@@
Let $H=\{h_{1},\ldots,h_{m}\}$ be a set of homogeneous polynomials in $R^{\prime}\smallsetminus K$ with $R^{\prime}=R[y]$, and put $\bm{H}:=(h_{1},\ldots,h_{m})\in(R^{\prime}\smallsetminus K)^{m}$. Assume that $R/\langle H^{\rm top}\rangle$ is Artinian, and that $\bm{H}^{\rm top}$ is cryptographic semi-regular (and hence $\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})<\infty$ by Theorem 3.1.1, see also Remark 3.1.2). Then we have the following:
@@
    $\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})\geq d_{\rm reg}(\langle H^{\rm top}\rangle_{R})-1$.
@@
    $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\leq\max\{{d}_{\rm reg}(\langle H^{\rm top}\rangle_{R}),\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})\}$ for any graded monomial ordering $\prec^{\prime}$ on $R^{\prime}$ given in Lemma 2.2.2.
Moreover, when we assume that $(H|_{y=1})^{h}=H$, the equality holds in the second inequality if $\prec^{\prime}$ is a DRL ordering and if $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [37, Section 4] for the original definition).
@@
Put $D:=\widetilde{d}_{\rm reg}(\langle H^{\rm top}\rangle_{R})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})$. Recall from the proof of Lemma 2.3.5 that we have the exact sequence (2.3.1). For any $d$ with $d\leq D-1$, we have $H_{0}(K_{\bullet}^{\prime})_{d}\neq 0$, and it follows from the cryptographic semi-regularity of $\bm{H}^{\rm top}$ that $H_{1}(K_{\bullet}^{\prime})_{d}=0$ by Proposition 2.1.11, whence $\dim_{K}(R^{\prime}/\langle H\rangle)_{d-1}<\dim_{K}(R^{\prime}/\langle H\rangle)_{d}$ for any such $d$. Therefore, the first assertion to be proved holds.
As in the proof of Lemma 2.3.5, it follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}^{\prime})_{d}=0$ for any $d$ with $d\geq D$. Thus, for such any $d$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle H\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle H\rangle)_{d}$.
@@
Here, we prove the second assertion $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\leq\max\{D,D^{\prime}\}$. If $D^{\prime}\geq D$, the assertion is a consequence of Lemma 2.3.5. Thus, it suffices to consider the case where $D^{\prime}=D-1$. In this case, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1=D$. However, $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$ is injective but not surjective, we cannot apply Lemma 2.2.2 for $d_{0}=D^{\prime}$, but apply it for $d_{0}=D^{\prime}+1=D$. Hence, $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)$ is upper-bounded by $D^{\prime}+1=D=\max\{D,D^{\prime}\}$, as desired.
Now, assuming that $(H|_{y=1})^{h}=H$, we show that the equality holds in the second inequality if $\prec^{\prime}$ is a DRL ordering and if $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic. Let $G_{H}$ be the reduced Gröbner basis of $\langle H\rangle_{R^{\prime}}$ with respect to $\prec^{\prime}$. Note that the elements of $\mathrm{LM}(G_{H})$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}(\langle H\rangle)\rangle=\langle\mathrm{LM}(G_{H})\rangle$.
@@
First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$. Put $d:=\mathrm{max.GB.deg}_{\prec^{\prime}}(H)$. Then, it follows from Macaulay’s basis theorem (cf. [29, Theorem 1.5.7]) that $R^{\prime}_{d}/\langle H\rangle_{d}$ has a $K$-linear basis of the form
$\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle H\rangle)\rangle=\langle\mathrm{LM}(G_{H})\rangle}\},$
@@
which is called the standard monomial basis. We set $r:=\dim_{K}R^{\prime}_{d}/\langle H\rangle_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$. In the following, we prove that $\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ is a $K$-linear basis of $R^{\prime}_{d+s}/\langle H\rangle_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R^{\prime}_{d}/\langle H\rangle_{d}=\dim_{K}R^{\prime}_{d+s}/\langle H\rangle_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
@@
By Remark 3.1.2 and Lemma 2.3.8 below, the multiplication-by-$y^{s}$ map from $R_{d}^{\prime}/\langle H\rangle_{d}$ to $R^{\prime}_{d+s}/\langle H\rangle_{d+s}$ is surjective. Therefore $B_{s}:=\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ generates $R^{\prime}_{d+s}/\langle H\rangle_{d+s}$. Suppose to contrary that $B_{s}$ is not a basis. In this case, there exists an $i$ such that $t_{i}y^{s}$ is divisible by $\mathrm{LM}(g)$ for some $g\in G_{H}$. Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}y^{s}=t_{i}y^{s}$. Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}(g)$, we can write $u_{i}y^{s^{\prime}}=\mathrm{LM}(g)$ for some $s^{\prime}\leq s$. (We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}(g)$.
@@
Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$. Then, by the weakly reverse lexicographicness, as ${\rm LM}(g)=u_{i}y^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}(G_{H})\rangle$. Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}(G_{H})\rangle$, which is a contradiction.
@@
Finally, we consider the remaining case, namely we have $D>D^{\prime}$, so that $\max\{D,D^{\prime}\}=D$. In this case, it follows from Lemma 2.3.8 below that $d\geq D$, whence $d=D$. ∎
@@
with $A^{\prime}:=R^{\prime}/\langle H\rangle$, and the case where $D^{\prime}=D-1$ (i.e., $D^{\prime}+1=D$) means that
@@
Let $H$, $\bm{H}$, and $\prec^{\prime}$ be as in Proposition 2.3.6. Assume that $R/\langle H^{\rm top}\rangle$ is Artinian, and that $\bm{H}^{\rm top}$ is cryptographic semi-regular. We also suppose that $\prec^{\prime}$ is a DRL ordering and that $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic. Then we have $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\geq d_{\rm reg}(\langle H^{\rm top}\rangle_{R})$ and $\mathrm{max.GB.deg}_{\prec}(H^{\rm top})=d_{\rm reg}(\langle H^{\rm top}\rangle_{R})$, where $\prec$ denotes the restriction of $\prec^{\prime}$ to $R$.